# On finite dimensional regular gradings

### v2

@@ Abstract.
Let $A$ be an associative algebra over an algebraically closed field $K$ of characteristic 0. A decomposition $A=A_{1}\oplus\cdots\oplus A_{r}$ of $A$ into a direct sum of $r$ vector subspaces is called a regular decomposition if, for every $n$ and every $1\leq i_{j}\leq r$, there exist $a_{i_{j}}\in A_{i_{j}}$ such that $a_{i_{1}}\cdots a_{i_{n}}\neq 0$, and moreover, for every $1\leq i,j\leq r$ there exists a constant $\beta(i,j)\in K^{*}$ such that $a_{i}a_{j}=\beta(i,j)a_{j}a_{i}$ for every $a_{i}\in A_{i}$, $a_{j}\in A_{j}$. We work with decompositions determined by gradings on $A$ by a finite abelian group $G$. In this case, the function $\beta\colon G\times G\to K^{*}$ ought to be a bicharacter. A regular decomposition is minimal whenever for every $g$, $h\in G$, the equalities $\beta(x,g)=\beta(x,h)$ for every $x\in G$ imply $g=h$. In this paper we describe the finite dimensional algebras $A$ (with unit) admitting a $G$-grading such that the corresponding regular decomposition is minimal. Moreover we compute the graded codimension sequence of these algebras assuming complete support. It turns out that the graded PI exponent of every finite dimensional $G$-graded algebra with regular grading such that its regular decomposition is minimal and it has complete support, coincides with its ordinary (ungraded) PI exponent. Finally we have shown that the regular decomposition of a finite dimensional regular $G$-grading AA is minimal if and only if $\exp(A)=|G|$.
@@ 1. Introduction
In this paper we give a description of the finite dimensional $G$-graded algebras whose decomposition is regular and minimal, and we give a complete description of these algebras with complete support, assuming $G$ a finite abelian group and the base field $K$ algebraically closed and of characteristic 0. In order to achieve these we use the graded version of the Wedderburn–Malcev decomposition of a finite dimensional algebra, and reduce the study to that of twisted group algebras. As a consequence of the above description, we also compute the graded codimension sequence of algebras with regular grading such that the regular decomposition is minimal and it has complete support. It turns out that in the last case, the graded PI exponent of $A$ coincides with the usual (ungraded) one. Eventually, we give a characterization of finite dimensional $G$-graded regular algebras with minimal decomposition in terms of its PI-exponent.
@@ 3. A motivation: regular $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$-gradings
4. Description of the structure of finite dimensional regular gradings with minimal regular decomposition and complete support
@@
[19, Theorem 2.2] Let $Q_{0}(\alpha)$ denote the subgroup of $Q$ consisting of all regular elements. The following conditions are equivalent:
@@
In what follows we shall use the following characterization of the PI-exponent of $A$, denoted by $\exp(A)$. Consider the Wedderburn-Malcev decomposition of $A=B_{1}\oplus B_{2}\oplus\cdots\oplus B_{k}+J(A)$. Then, $\exp(A)$ is equal to the maximal value of the sum of the dimensions
@@
Let $B$ be a finite dimensional $G$-graded algebra with regular grading. We say $B$ has complete support if any twisted group algebra contained in $B/J(B)$ is isomorphic to $K^{\alpha}G$, where $\alpha\in H^{2}(G,K^{\ast})$.
@@
    There exists $1\leq i\leq k$ such that $D_{i}$ is isomorphic to $K^{\alpha}G$, where $\alpha\in H^{2}(G,K^{\ast})$. If $J(A)\neq 0$ and the regular decomposition of $A$ is minimal, then there exists $1\leq j\leq k$ such that $D_{j}J(A)\neq 0$.
@@
    If the regular decomposition of $A$ is minimal and $A$ has complete support, then
@@
It follows from the graded Wedderburn-Malcev theorem [9, Lemma 2.2] that the Jacobson radical $J=J(A)$ is graded with respect to the $G$-grading and there exists a semisimple subalgebra $D$ that is homogeneous in the $G$-grading such that $A=D+J$ as a direct sum of vector subspaces. Moreover, $D$ can be decomposed as the direct sum $D=D_{1}\oplus D_{2}\oplus\cdots\oplus D_{k}$ of graded two-sided ideals of $D$ and each $D_{j}$ is a $G$-graded simple algebra. We also note that by construction $D_{i}D_{j}=0$ for $i\neq j$. By Proposition 24 it follows that $D_{i}\cong K^{\alpha_{i}}H_{i}$ (as $G$-graded algebras), where $H_{i}=\operatorname{Supp}(D_{i})$ is a subgroup of $G$, for all $1\leq i\leq k$. Of course $J$ cannot have a regular grading because it is nilpotent. Therefore, $D$ must have a $G$-graded regular grading, and since $K^{\alpha_{i}}H_{i}$ it is not a $G$-graded regular algebra if $H_{i}$ is a proper subgroup of $G$, we conclude there exists $1\leq i\leq k$ such that $D_{i}\cong K^{\alpha_{i}}G$.
Suppose the regular decomposition of $A$ is minimal, then by Theorem [1] we have $\exp(A)=|G|$. Therefore since $J(A)\neq 0$, by definition of the PI-exponent we conclude there exists $D_{j}\cong K^{\alpha_{j}}G$ such that $D_{j}J(A)\neq 0$, and (1) follows.
@@
Lemma 28.
Let $A$ be a finite dimensional algebra with regular $G$-grading and bicharacter $\beta$. Assume the regular decomposition of $A$ is minimal and $J(A)\neq 0$. Write $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$ where $D_{i}$ is a twisted group algebra. Then, for each $D_{i}\cong K^{\alpha_{i}}G$ it follows that $D_{i}J(A)\neq 0$.
Proof.
First suppose $k=1$. Then, since the unit of $A$ is in $D_{1}$, if $D_{1}J(A)=0$ we would have $J(A)=0$ and that is a contradiction. Now suppose $k>1$ and take some $1\leq s\leq k$ such that $D_{s}\cong K^{\alpha_{s}}G$ and $D_{s}J(A)=0$. Of course, $\alpha_{s}$ is a cocycle that induces $\beta$. We set $C:=D_{s}\oplus J(A)$. Since $D_{s}J(A)=0$ it follows that $C$ is a $G$-graded algebra (without unit) which is $\beta$-commutative. On the other hand, the fact that $D_{s}$ has a regular grading assures us that $C$ has one as well, by Remark 27. Further, the bicharacter of $C$ is $\beta$. In particular the regular decomposition of $C$ is minimal. By Theorem 10 we have $\exp(C)=|G|$, and so, $D_{s}J(A)\neq 0$, that is a contradiction. We conclude $D_{s}J(A)\neq 0$. ∎
@@
Let $A$ be an algebra as in Proposition 26. Assume that the regular decomposition of $A$ is minimal and $A$ has complete support. Without loss of generality we can take $A=K^{\alpha_{1}}G\oplus\cdots\oplus K^{\alpha_{k}}G+J(A)$, with $\alpha_{1}=\cdots=\alpha_{k}=\alpha$.
@@
    Since the regular decomposition of $A$ is minimal, by Theorem 10, it follows that $\det M^{A}\neq 0$ and $\exp(A)=|G|$. It is worth noting the cocycle $\alpha$ cannot be trivial, otherwise, $\beta\equiv 1$ and $\det M^{A}=0$.
@@
    By Lemma 20 we get the twisted group algebra $K^{\alpha_{j}}G$ is central simple. Consequently, the decomposition $A=K^{\alpha_{1}}G\oplus\cdots\oplus K^{\alpha_{k}}G+J(A)$ is the ordinary Wedderburn-Malcev decomposition since all $K^{\alpha_{j}}G$ are simple. Hence, if $J(A)\neq 0$ and there exist $1\leq s\neq j\leq k$ such that
    $(K^{\alpha_{j}}G)J(A)(K^{\alpha_{s}}G)\neq 0,$
    then
    $\dim(K^{\alpha_{j}}G)+\dim(K^{\alpha_{s}}G)\leq\exp(A),$
    that is a contradiction because $\exp(A)=|G|$ and $\dim K^{\alpha_{j}}G=\dim K^{\alpha_{s}}G=|G|$. It follows $(K^{\alpha_{i}}G)J(A)(K^{\alpha_{j}}G)=0$ for any $1\leq i,j\leq k$.
@@
Let $A$ be a finite dimensional $G$-graded algebra with regular grading. Then $A\cong K^{\alpha}G+J(A)$, if and only if $A_{0}$ is a local algebra.
@@
Conversely, suppose that $A_{0}=K+J(A_{0})$. By Proposition 26, there exist graded simple subalgebras $D_{1}$, …, $D_{k}$ of $A$ such that $D_{i}\cong K^{\alpha}H_{i}$, $\alpha_{i}\in H^{2}(G,K^{\ast})$, for each $1\leq i\leq k$. Also, again by Proposition 26 there exists $D_{i}\cong K^{\alpha_{i}}G$. Assume $k>1$ and let $\{d_{g}^{(1)}\mid g\in H_{1}\}$ and $\{d_{g}^{(2)}\mid g\in H_{2}\}$ be bases of $D_{1}$ and $D_{2}$, respectively, with $d_{g}^{(i)}d_{h}^{(i)}=\alpha(g,h)d_{g+h}^{(i)}$ for $i\in\{1,2\}$ and $g$, $h\in H_{i}$. It is clear that $d_{0}^{(2)}\in K$ because $J(D_{2})=0$ and $(D_{2})_{0}\subseteq A_{0}$. By similar consideration, $d_{0}^{(1)}\in K$, implying the existence of $\lambda\in K$ such that $d_{0}^{(1)}=\lambda d_{0}^{(2)}$, that is a contradiction. We conclude that $i=1$, thus $A\cong K^{\alpha}G+J(A)$. ∎
Finite dimensional regular graded algebras whose neutral component is local do not necessarily need to be isomorphic to a twisted group algebra, even if the regular decomposition is minimal; this is well illustrated by the following two examples.
@@
$U_{T}:=\operatorname{span}_{K}\{e_{1,1}+e_{2,2},e_{1,2}\}=\Big\{\begin{pmatrix}a&b\\ 0&b\end{pmatrix}\mid a,b\in K\Big\}.$
Then, by Remark 15, the $\{0\}$-graded algebra $U_{T}=(U_{T})_{0}$ is a regular algebra with minimal regular decomposition, and it is a local algebra because $U_{T}=K(e_{1,1}+e_{1,2})+J(U_{T})$, where $J(U_{T})=Ke_{1,2}$. However, $U_{T}$ is not isomorphic to a twisted group algebra, since by [25, Theorem 2], a twisted group algebra (over a finite group and an algebraically closed field of characteristic 0) is semisimple, while $J(U_{T})\neq 0$.
@@
Let $A$ be a finite dimensional algebra with regular $G$-grading and bicharacter $\beta$. Assume $A_{0}$ is a local algebra. Then, if $J(A_{0})\neq 0$, we have
@@
Let $A$ be a finite dimensional algebra with regular $G$-grading and bicharacter $\beta$. Assume $A_{0}$ is a local algebra. Then, if $J(A_{0})\neq 0$, there exist $\mathscr{U}$ a commutative local subalgebra of $A$ and $\alpha\in H^{2}(G,K^{\ast})$ be a cocycle that induces $\beta$ such that $A\cong K^{\alpha}G\otimes\mathscr{U}$ as $G$-graded algebras (here we consider $\mathscr{U}$ with the trivial grading). Moreover, if $J(A_{0})=0$, then $A\cong K^{\alpha}G$.
@@
Now we consider regular gradings with minimal decompositions where the component $A_{0}$ is not necessarily local. By the Wedderburn-Malcev theorem we write $A_{0}=E_{1}\oplus\cdots\oplus E_{l}+J(A_{0})$, where $E_{i}\cong K$, for every $1\leq i\leq l$. Denote by $1_{E_{i}}$ the identity of $E_{i}$, $1\leq i\leq l$, then we have $1=1_{E_{1}}+\cdots+1_{E_{l}}$.
@@
Let $A$ be a finite dimensional $G$-graded algebra with regular grading and bicharacter $\beta$. Assume that the regular decomposition of $A$ is minimal and it has complete support. Let $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$, where $D_{i}\cong K^{\alpha}G$, for every $1\leq i\leq k$, and $\alpha$ induces $\beta$. For each $1\leq i\leq k$, let $\mathcal{D}_{i}:=\{d_{g}^{(i)}\mid g\in G\}$ be a basis of $D_{i}$ satisfying
$d_{g}^{(i)}d_{h}^{(i)}=\alpha_{i}(g,h)d_{g+h}^{(i)},\quad\text{for any}\quad g,h\in G.$
@@
We know that for each $1\leq i\leq k$, $D_{i}$ is a $G$-graded regular algebra with bicharacter $\beta$ induced by $\alpha$. It follows by [25, Theorem 2] that $J(D_{i})=0$, and in particular $J((D_{i})_{0})=0$. Therefore by Corollary 30 we conclude $(D_{i})_{0}\cong K$. On the other hand, $(D_{i})_{0}\subseteq A_{0}$, then $(D_{i})_{0}$ must be equal to $E_{r_{i}}$, for some $1\leq r_{i}\leq l$. In particular, $d_{0}^{(i)}\in E_{r_{i}}$.
@@
Finally, we will show that $E_{r_{i}}J(A_{0})\neq 0$ for every $1\leq i\leq k$. We already know by Lemma 28 that $D_{i}J(A)\neq 0$, for each $1\leq i\leq k$. Let $1\leq i\leq k$ and take $t\in D_{i}$ and $a\in J(A)_{g}$ (for some $g\in G$) such that $ta\neq 0$. If $1_{E_{r_{i}}}a=0$, then $d_{h}^{(i)}a=0$, for any $h\in G$ and so, $ta=0$, that is a contradiction. Thus, $1_{E_{r_{i}}}a\neq 0$. In particular, $1_{E_{r_{i}}}(d_{g}^{(i)})^{-1}a=(d_{g}^{(i)})^{-1}a\neq 0$, i.e, $E_{r_{i}}J(A_{0})\neq 0$. ∎
@@
Let $A$ be a finite dimensional $G$-graded algebra with regular grading and bicharacter $\beta$. Assume that the regular decomposition of $A$ is minimal and $A$ has complete support. Suppose $J(A_{0})\neq 0$, then, considering $J(A_{0})$ with the trivial grading we have
@@
Let $A$ be a finite dimensional $G$-graded algebra with regular grading and bicharacter $\beta$. Assume that the regular decomposition of $A$ is minimal and $A$ has complete support. Suppose that $J(A_{0})\neq 0$. Then, there exists $\mathscr{V}$ a commutative algebra such that the following conditions are satisfied

### v3

@@ Abstract.
Let $A$ be an associative algebra over an algebraically closed field $K$ of characteristic 0. A decomposition $A=A_{1}\oplus\cdots\oplus A_{r}$ of $A$ into a direct sum of $r$ vector subspaces is called a regular decomposition if, for every $n$ and every $1\leq i_{j}\leq r$, there exist $a_{i_{j}}\in A_{i_{j}}$ such that $a_{i_{1}}\cdots a_{i_{n}}\neq 0$, and moreover, for every $1\leq i,j\leq r$ there exists a constant $\beta(i,j)\in K^{*}$ such that $a_{i}a_{j}=\beta(i,j)a_{j}a_{i}$ for every $a_{i}\in A_{i}$, $a_{j}\in A_{j}$. We work with decompositions determined by gradings on $A$ by a finite abelian group $G$. In this case, the function $\beta\colon G\times G\to K^{*}$ ought to be a bicharacter. A regular decomposition is minimal whenever for every $g$, $h\in G$, the equalities $\beta(x,g)=\beta(x,h)$ for every $x\in G$ imply $g=h$. In this paper we describe completely the structure of the finite dimensional algebras $A$ (with unit) admitting a $G$-regular grading. Moreover, we compute the graded codimension sequence for a class of such algebras assuming complete support and minimal regular decomposition. It turns out that, for these algebras, the graded PI-exponent coincides with the ordinary (ungraded) PI-exponent. Finally, we show that the regular decomposition of a finite-dimensional algebra $A$ with a regular $G$-grading is minimal if and only if $\exp(A)=|G|$.
@@ 1. Introduction
In this paper we give a complete description of the finite dimensional $G$-graded regular algebras assuming $G$ a finite abelian group and the base field $K$ algebraically closed and of characteristic 0. In order to achieve these we use the graded version of the Wedderburn–Malcev decomposition of a finite dimensional algebra, and reduce the study to that of twisted group algebras. As a consequence of the description above, we also compute the graded codimension sequence of algebras $A$ with a regular grading whose regular decomposition is minimal, has complete support, and whose graded simple components have nonzero product with the Jacobson radical. It turns out that, in this case, the graded PI-exponent of $A$ coincides with the usual (ungraded) one. Finally, we provide a characterization of finite-dimensional $G$-graded regular algebras with minimal decomposition in terms of their PI-exponent.
@@ 3. A motivation: regular $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$-gradings
4. Description of the structure of finite dimensional regular gradings with complete support
@@
[20, Theorem 2.2] Let $Q_{0}(\alpha)$ denote the subgroup of $Q$ consisting of all regular elements. The following conditions are equivalent:
@@
Recall the characterization of the PI-exponent of $A$, denoted by $\exp(A)$. Consider the Wedderburn-Malcev decomposition of $A=B_{1}\oplus B_{2}\oplus\cdots\oplus B_{k}+J(A)$. Then, $\exp(A)$ is equal to the maximal value of the sum of the dimensions
@@
We observe that, in the definition of the PI-exponent the dimensions of all simple subalgebras of $R$ are also taken into account.
@@
Let $B$ be a finite dimensional $G$-graded algebra with regular grading. We say $B$ has complete support if any twisted group algebra contained in $B/J(B)$ is isomorphic to $K^{\alpha}G$, for some $\alpha\in H^{2}(G,K^{\ast})$.
@@
    There exists $1\leq i\leq k$ such that $D_{i}$ is isomorphic to $K^{\alpha}G$, where $\alpha\in H^{2}(G,K^{\ast})$.
@@
    If $A$ has complete support, then
@@
It follows from the graded Wedderburn-Malcev theorem [5, Lemma 2.2] that the Jacobson radical $J=J(A)$ is graded with respect to the $G$-grading and there exists a semisimple subalgebra $D$ that is homogeneous in the $G$-grading such that $A=D+J$ as a direct sum of vector subspaces. Moreover, $D$ can be decomposed as the direct sum $D=D_{1}\oplus D_{2}\oplus\cdots\oplus D_{k}$ of graded two-sided ideals of $D$ and each $D_{j}$ is a $G$-graded simple algebra. We also note that by construction $D_{i}D_{j}=0$ for $i\neq j$. By Proposition 24 it follows that $D_{i}\cong K^{\alpha_{i}}H_{i}$ (as $G$-graded algebras), where $H_{i}=\operatorname{Supp}(D_{i})$ is a subgroup of $G$, for all $1\leq i\leq k$. Of course $J$ cannot have a regular grading because it is nilpotent. Therefore, $D$ must have a $G$-graded regular grading, and since $K^{\alpha_{i}}H_{i}$ it is not a $G$-graded regular algebra if $H_{i}$ is a proper subgroup of $G$, we conclude there exists $1\leq i\leq k$ such that $D_{i}\cong K^{\alpha_{i}}G$.
@@
Let $A$ be a finite dimensional algebra with regular $G$-grading and bicharacter $\beta$. By Proposition 26 we can write $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$, where $D_{1}$,…, $D_{k}$ are twisted group algebras with at least one isomorphic to $K^{\alpha}G$, $\alpha\in H^{2}(G,K^{\ast})$. If $k=1$, since we are assuming $A$ has a unit it follows that $D_{1}J(A)\neq 0$. Suppose $k>1$ and assume $D_{1}J(A)=J(A)D_{1}=0$. Then, of course, in this case $A$ must be isomorphic to the algebra
$\begin{pmatrix}B&0\\ 0&D_{1}\end{pmatrix}$
where $B=D_{2}\oplus\cdots\oplus D_{k}+J(B)$, $J(B)=J(A)$. If $D_{1}\cong K^{\alpha}G$, then $B$ may be any $\beta$-commutative algebra, with or without a regular $G$-grading. If $D_{1}\cong K^{\alpha_{i}}H_{i}$, where $H_{i}$ is a proper subgroup of $G$, and $\alpha_{i}\in H^{2}(H_{i},K^{\ast})$, then we look at $D_{2}J(A)$ and this process can be continued. The next example illustrates a case in which this occurs.
Example 28.
In $M_{6}(K)$, consider the following matrices
$L=\begin{pmatrix}D&0&0\\ 0&D&0\\ 0&0&D\end{pmatrix},\quad M:=\begin{pmatrix}0&N&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad J=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&N\end{pmatrix}$
where
$D=\operatorname{diag}(1,-1),\quad\text{and}\quad N=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$.
These matrices satisfies the following relations
$LM=-ML,\quad M^{2}=0,\quad JM=MJ=0,\quad LJ=-JL.$
Moreover $N^{2}=I_{2}$, $L^{2}=I_{6}$, where $I_{2}$ and $I_{6}$ are the identity matrices in $2\times 2$ and $6\times 6$ respectively, $D^{n}=\operatorname{diag(1,(-1)^{n})}$ for all $n\in\mathbb{N}$, and
$LJ=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&P\end{pmatrix},\quad\text{where}\quad P=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$.
Thus, $(LJ)^{2}=\operatorname{diag(0,0,I_{2})}$. Therefore if $\mathcal{A}$ is the subalgebra of $M_{6}(K)$ generated by $L$, $M$ and $J$, it can be seen that $\{I,J,J^{2},L,LJ,LJ^{2},M,LM\}$ is a basis of $\mathcal{A}$. Now, it is not difficult to see that $\mathcal{A}$ is a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded algebra if we put
$\mathcal{A}_{(0,0)}:=\operatorname{span}_{K}\{I_{6},J^{2}\},\quad\mathcal{A}_{(0,1)}:=\operatorname{span}_{K}\{L,LJ^{2}\},\quad\mathcal{A}_{(1,0)}:=\operatorname{span}_{K}\{M,J\}$
and $\mathcal{A}_{(1,1)}:=\operatorname{span}_{K}\{LM,LJ\}$.
Now, notice the following
$(I_{6})(L)(J)(LJ)=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&-I_{2}\end{pmatrix}$
which is non-nilpotent. Therefore, by the above relations, it follows that $\mathcal{A}$ is a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded regular algebra whose the decomposition matrix is given by
$M^{\mathcal{A}}=\begin{pmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{pmatrix}$.
Now, we will show that $J(\mathcal{A})=\operatorname{span}_{K}\{M,ML\}$. Consider $\mathcal{I}$ the ideal generated by $M$. By construction $\mathcal{I}=\operatorname{span}_{K}\{M,ML\}$, then it can be seen that
$\mathcal{A}/\mathcal{I}=\mathcal{B}$
where $\mathcal{B}$ is isomorphic to the subalgebra $\operatorname{span}_{K}\{I_{6},J,J^{2},L,LJ,LJ^{2}\}$. It can be easily seen that $\operatorname{span}_{K}\{I_{2},P,N,D\}$ is $\langle D,N\rangle$, the subalgebra generated by $D$ and $N$, and it is isomorphic to $M_{2}(K)$. Thus, since $LJ^{2}=\operatorname{diag}(0,0,D)$, the subalgebra
$\langle J,LJ^{2}\rangle=\operatorname{span}_{K}\{J,J^{2},LJ,LJ^{2}\}$
is isomorphic to $M_{2}(K)$. Now, defining $\mathscr{E}:=\operatorname{diag}(D,D,0)$, it follows that $\mathscr{E}^{2}=\operatorname{diag}(I_{2},I_{2},0)=I_{4}$, the $4\times 4$ identity matrix, and we have
$L=\mathscr{E}+LJ^{2},\quad\mathscr{E}(LJ^{2})=0,\quad\text{and}\quad I_{6}=I_{4}+(LJ^{2})^{2}.$
On the other hand, if $\mathcal{C}:=\operatorname{span}_{K}\{\mathscr{E},I_{4}\}$, then $\mathcal{C}$ is a subalgebra and
$\mathcal{C}\cong K[x]/(x^{2}-1)\cong K\oplus K$
where the last direct sum is as algebras. Therefore, it follows that
$\mathcal{B}\cong K\oplus K\oplus M_{2}(K)\quad(\text{direct sum as algebras}).$
We conclude $J(\mathcal{A}/\mathcal{I})=0$. Therefore $\mathcal{I}=J(\mathcal{A})$. In this case, since $\langle J,LJ^{2}\rangle J(A)=0$ the algebra $\mathcal{A}$ is isomorphic to $\begin{pmatrix}B&0\\ 0&M_{2}(K)\end{pmatrix}$, where $B=K\oplus K+\operatorname{span}_{K}\{M,ML\}$.
With the illustration provided by the Example 28, and by the discussion preceding it, we may henceforth assume, without loss of generality, that $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$ is a regular $G$-graded algebra with bicharacter $\beta$ and that if $J(A)\neq 0$, then $D_{i}J(A)\neq 0$ for all $1\leq i\leq k$. In Section  6, it will become clear from the analysis how to deal with the case in which $A$ is only $\beta$-commutative.
@@
    If the regular decomposition of $A$ is minimal, by Theorem 10, it follows that $\det M^{A}\neq 0$ and $\exp(A)=|G|$. It is worth noting the cocycle $\alpha$ cannot be trivial, otherwise, $\beta\equiv 1$ and $\det M^{A}=0$.
@@
    Since $A$ is in particular $\beta$-commutative, given $i\neq j$, because $D_{i}D_{j}=0$, it follows that $D_{i}J(A)D_{j}=0$.
@@
$A\cong K^{\alpha}G+J(A)$, if and only if $A_{0}$ is a local algebra.
@@
Conversely, suppose that $A_{0}=K+J(A_{0})$. By Proposition 26, there exist graded simple subalgebras $D_{1}$, …, $D_{k}$ of $A$ such that $D_{i}\cong K^{\alpha}H_{i}$, $\alpha_{i}\in H^{2}(G,K^{\ast})$, for each $1\leq i\leq k$. Also, again by Proposition 26 there exists $D_{i}\cong K^{\alpha_{i}}G$. Assume $k>1$ and let $\{d_{g}^{(1)}\mid g\in H_{1}\}$ and $\{d_{g}^{(2)}\mid g\in H_{2}\}$ be bases of $D_{1}$ and $D_{2}$, respectively, with $d_{g}^{(i)}d_{h}^{(i)}=\alpha(g,h)d_{g+h}^{(i)}$ for $i\in\{1,2\}$ and $g$, $h\in H_{i}$. It is clear that $d_{0}^{(1)}$, $d_{0}^{(2)}\in K$ because $(D_{i})_{0}\cong K$, $i\in\{1,2\}$. It implies the existence of $\lambda\in K$ such that $d_{0}^{(1)}=\lambda d_{0}^{(2)}$, that is a contradiction. We conclude that $i=1$, thus $A\cong K^{\alpha}G+J(A)$. ∎
Finite-dimensional regular graded algebras whose neutral component is local do not necessarily have to be isomorphic to a twisted group algebra, even when the regular decomposition is minimal. This is well illustrated by the following two examples, which rely on the fact that a twisted group algebra (over an algebraically closed field of characteristic 0) is semisimple (see [19, Theorem 3.3.6]).
@@
$U_{T}:=\operatorname{span}_{K}\{e_{1,1}+e_{2,2},e_{1,2}\}=\Big\{\begin{pmatrix}a&b\\ 0&a\end{pmatrix}\mid a,b\in K\Big\}.$
Then, by Remark 15, the $\{0\}$-graded algebra $U_{T}=(U_{T})_{0}$ is a regular algebra with minimal regular decomposition, and it is a local algebra because $U_{T}=K(e_{1,1}+e_{1,2})+J(U_{T})$, where $J(U_{T})=Ke_{1,2}$. Because $J(U_{T})\neq 0$, $U_{T}$ cannot be isomorphic to a twisted group algebra.
@@
Assume $A_{0}$ is a local algebra. Then, if $J(A_{0})\neq 0$, we have
@@
Assume $A_{0}$ is a local algebra. Then, if $J(A_{0})\neq 0$, there exist $\mathscr{U}$ a commutative local subalgebra of $A$ and $\alpha\in H^{2}(G,K^{\ast})$ be a cocycle that induces $\beta$ such that $A\cong K^{\alpha}G\otimes\mathscr{U}$ as $G$-graded algebras (here we consider $\mathscr{U}$ with the trivial grading). Moreover, if $J(A_{0})=0$, then $A\cong K^{\alpha}G$.
@@
Now we consider the case where $A_{0}$ is not necessarily local. By the Wedderburn-Malcev theorem we write $A_{0}=E_{1}\oplus\cdots\oplus E_{l}+J(A_{0})$, where $E_{i}\cong K$, for every $1\leq i\leq l$. Denote by $1_{E_{i}}$ the identity of $E_{i}$, $1\leq i\leq l$, then we have $1=1_{E_{1}}+\cdots+1_{E_{l}}$.
@@
Assume that $A$ has complete support. Write $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$, where $D_{i}\cong K^{\alpha}G$, for every $1\leq i\leq k$, and $\alpha$ induces $\beta$. For each $1\leq i\leq k$, let $\mathcal{D}_{i}:=\{d_{g}^{(i)}\mid g\in G\}$ be a basis of $D_{i}$ satisfying
$d_{g}^{(i)}d_{h}^{(i)}=\alpha(g,h)d_{g+h}^{(i)},\quad\text{for any}\quad g,h\in G.$
@@
For each $1\leq i\leq k$, $D_{i}$ is a twisted group algebra isomorphic to $K^{\alpha}G$, in particular $D_{i}$ is a $G$-graded regular algebra with bicharacter $\beta$. We already know that $(D_{i})_{0}\cong K$. On the other hand, $(D_{i})_{0}\subseteq A_{0}$, then $(D_{i})_{0}$ must be equal to $E_{r_{i}}$, for some $1\leq r_{i}\leq l$. In particular, $d_{0}^{(i)}\in E_{r_{i}}$.
@@
Finally, we will show that $E_{r_{i}}J(A_{0})\neq 0$ for every $1\leq i\leq k$. Recall that, by hypothesis on $A$, $D_{i}J(A)\neq 0$, for each $1\leq i\leq k$. Let $1\leq i\leq k$ and take $t\in D_{i}$ and $a\in J(A)_{g}$ (for some $g\in G$) such that $ta\neq 0$. If $1_{E_{r_{i}}}a=0$, then $d_{h}^{(i)}a=0$, for any $h\in G$ and so, $ta=0$, that is a contradiction. Thus, $1_{E_{r_{i}}}a\neq 0$. In particular, $1_{E_{r_{i}}}(d_{g}^{(i)})^{-1}a=(d_{g}^{(i)})^{-1}a\neq 0$, i.e, $E_{r_{i}}J(A_{0})\neq 0$. ∎
@@
Assume that $A$ has complete support. Suppose $J(A_{0})\neq 0$, then, considering $J(A_{0})$ with the trivial grading we have
@@
Assume that $A$ has complete support. Suppose that $J(A_{0})\neq 0$. Then, there exists $\mathscr{V}$ a commutative algebra such that the following conditions are satisfied